OpenStudy (anonymous):

can (x-7)/(7-x) be simplified

OpenStudy (tyteen4a03):

Yes. (7-x) == -(x-7).

OpenStudy (anonymous):

That appears to be the simplest form possible

OpenStudy (anonymous):

-x+7 and 7-x aren't the same are they?

OpenStudy (kirbykirby):

(7-x) = -(x-7) Thus, (x-7)/(7-x) = (x-7)/-(x-7) = -1

OpenStudy (kirbykirby):

-7+x = 7-x @jbovey

OpenStudy (kirbykirby):

To be more accurate, you should add the restriction that x can't be equal to 7

OpenStudy (anonymous):

But -(x-7) simplifies to -x+7 not 7-x @kirbykirby

OpenStudy (kirbykirby):

But the idea is to simplify the whole fraction. By "de-simplifying" it to -(x-7), you can cancel out the (x-7) terms and thus end up with -1, which is a lot simpler than (x-7)/(7-x)

OpenStudy (kirbykirby):

This is common to do in math. You often factor the numerator because you can find something that looks like it in the denominator that you can cancel out.

OpenStudy (kirbykirby):

For example, if you have (x^2-1)/(x-1), you would say = [(x-1)(x+1)]/(x-1) = (x+1) @jbovey

OpenStudy (anonymous):

I see what u mean now, thanks for the explanation bud. Cheers @kirbykirby

OpenStudy (kirbykirby):

Np :) This factoring and cancelling idea will become useful when you do limits.